# Learning Macroeconomic Policies based on Microfoundations: A Dynamic Stackelberg Mean Field Game Approach

### v2

@@ Abstract
The Lucas critique emphasizes the importance of considering the impact of policy changes on the expectations of micro-level agents in macroeconomic policymaking. However, the inherently self-interested nature of large-scale micro-agents, who pursue long-term benefits, complicates the formulation of optimal macroeconomic policies. This paper proposes a novel general framework named Dynamic Stackelberg Mean Field Games (Dynamic SMFG) to model such policymaking within sequential decision-making processes, with the government as the leader and households as dynamic followers. Dynamic SMFGs capture the dynamic interactions among large-scale households and their response to macroeconomic policy changes. To solve dynamic SMFGs, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which leverages the population distribution of followers to represent high-dimensional joint state and action spaces. In experiments, our method surpasses macroeconomic policies in the real world, existing AI-based and economic methods. It allows the leader to approach the social optimum with the highest performance, while large-scale followers converge toward their best response to the leader’s policy. Besides, we demonstrate that our approach retains effectiveness even when some households do not adopt the SMFG policy. In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policy-making issues.
@@ 1 Introduction
Macroeconomic policymaking is fundamental to the sustained development of an economy [52, 46]. Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending [49]. Therefore, accurately modeling and solving optimal macroeconomic policies and simulating the effects of their implementation are critical issues. According to the Lucas critique [39], predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behavior—termed "microfoundations"—to forecast personal responses to policy changes and aggregate these into macroeconomic effects. Nonetheless, several challenges persist: dynamic interactions among micro-agents, between micro-agents and macro-authorities, the inherently selfish behavior of agents focused on optimizing long-term returns, and the curse of dimensionality due to the large-scale nature of agents such as households and firms.
This paper introduces a novel framework termed Dynamic Stackelberg Mean Field Game (Dynamic SMFG) designed to address complex macroeconomic policy-making problems. Within this dynamic setting, the government, serving as the leader, adjusts policies iteratively, while a large population of households, acting as followers, respond accordingly. This process involves sequential decision-making, with both the leader and the followers aiming to optimize their cumulative long-term benefits. Given the leader’s policy, the SMFG reduces to a mean field game, which is widely used for modeling interactions among large-scale homogenous agents [33]. In Dynamic SMFGs, followers aim to maximize their long-term benefits by providing the best response to the leader’s policy, while the leader’s goal is to find an optimal policy that effectively guides the massive followers toward achieving comprehensive macroeconomic objectives. Existing approaches often simplify Dynamic SMFGs to single-step decisions [45] or assume myopic followers [63], which do not adequately capture the complexities of macroeconomic policy-making. Therefore, we propose a data-driven reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), to solve Dynamic SMFGs without requiring transition information or prior knowledge. The SMFRL algorithm utilizes a mean field approximation [28] to reduce the high-dimensional joint state and actions of followers to a manageable population distribution, allowing for the effective learning of a central critic Q function based on the overall followers’ distribution.
Experiments conducted on TaxAI [41], an environment designed for studying optimal tax policies, first validate the performance of SMFRL in solving Dynamic SMFGs. The results indicate that SMFRL enables followers to closely achieve their best response while allowing the leader to approach the social optimum with the highest performance compared with baselines. Furthermore, the paper compares the Dynamic SMFG method with macroeconomic policies such as AI Economist [62], the Saez tax [50], the 2022 U.S. Federal tax, and free market policy [5], demonstrating the superiority of the Dynamic SMFG framework in macroeconomic policymaking. Considering real-world applications, our experiments show that even if a portion of households do not adopt the follower policy learned by the SMFG method, it does not adversely affect the utility, wealth, and income of households that do adopt the SMFG policy, which are in fact higher. However, the GDP of the economy may decrease with an increase in the number of households not supporting the SMFG policy, as they typically contribute less labor. In summary, the contributions of this paper are threefold:
1. It innovatively proposes a general framework based on Dynamic Stackelberg Mean Field Games to model macroeconomic policy-making, considering key issues such as microfoundations, dynamic interactions and feedback, rational expectations, and randomness.
2. It introduces a data-driven Stackelberg Mean Field Reinforcement Learning algorithm to solve dynamic SMFGs without environmental transition information and prior knowledge, validated experimentally for its effectiveness.
3. It empirically demonstrates the superiority of the SMFG method in modeling and solving optimal macroeconomic policy problems and discusses its robustness in real-world scenarios.
The source code for the SMFG method is available in an anonymous GitHub repository  [^1] https://anonymous.4open.science/r/SMFG_macro_7740.
@@ 2.1 Economic Methods
In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods. Classic theoretical models, such as the IS-LM model [24, 22], are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations. The aggregate demand-aggregate supply (AD-AS) model [19, 34] focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics. The Solow growth model [10] provides a framework for economic growth but does not consider market imperfections and externalities. New Keynesian models [9, 21] emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions. DSGE models [1, 55], built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections. Empirical analysis [47, 59] primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy. Econometric methods [30, 15], such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
2.2 AI for Economics
@@
Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems [56]. In terms of macroeconomics, AI Economist [62] employs curriculum learning on tax policy design, [58] studies the collaboration between central and local governments under COVID-19, as well as monetary policy [26, 12], international trade [51], market pricing [14]. However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies. [31] studies democratic AI by human voting models, [60] examines the impact of social sanction rules on labor division. However, these works are based on simplified settings and remain a gap in real-world economic policy. In terms of microeconomics, [53, 48, 3] study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium [32, 25], multiple equilibria under real-business-cycle [13], the emergence of barter behavior [29], optimal asset allocation and savings strategies [44]. In summary, research on macroeconomic problems still lacks a bridge framework of macro policy and microfoundations, which has drawn our attention to the Stackelberg mean field game.
Figure 1: The mapping from optimal macroeconomic policy problems (left) to the Dynamic Stackelberg Mean Field Games (right). The left represents the reality where the macro-level government implements fiscal or monetary policies to optimize GDP, welfare, or equity, and the micro-level consists of large-scale households engaging in economic activities. In dynamic SMFGs, the government acts as the leader, and households as followers, respond to the leader’s policy.
2.3 Stackelberg Mean Field Game
Stackelberg Mean Field Games (SMFG) model interactions between the leader and numerous homogeneous followers, employing both model-based and model-free approaches. Model-based methods, such as those presented in [23], treat SMFGs as minimax optimization problems, streamlining the leader’s strategy to a single optimal action rather than a comprehensive policy. Studies by [7, 20, 16, 8] utilize forward-backward stochastic functional differential equations to determine the followers’ mean field equilibrium before computing the leader’s optimal strategy. This includes addressing terminal state constraints [20] and simplifying scenarios via linear-quadratic formulations [6, 37, 43, 18, 27]. Such methods, though theoretically robust, often rely on simplifying real-world complexities and require extensive knowledge of environmental transitions. Extending SMFG to model epidemic dynamics [4] showcases its broader applicability in public health management. Conversely, model-free approaches leverage complex, data-rich environments without the need for prior environmental knowledge. [11] employs a modified deep backward stochastic differential equation framework to establish followers’ equilibrium, despite constraints on leader decisions. [45] investigates applications in privacy protection through a single-step decision model. [42] delves into locally optimal defensive strategies given predefined attacker trajectories. [35] learns followers’ transition functions from empirical data, and facilitates SMFG solutions through the empirical Fokker-Planck equation.
However, in real-world settings, the sequential decision-making between a leader and numerous followers to optimize long-term gains is common but inadequately studied. This paper proposes a data-driven reinforcement learning algorithm for solving the aforementioned problem, termed the Dynamic Stackelberg Mean Field Game, without reliance on environmental transition information or prior knowledge.
@@ 3 Stackelberg Mean Field Game Modeling
3.1 Macroeconomic Policy-Making
Macroeconomic policy-making involves formulating effective policies to regulate and guide an economy toward achieving sustainable growth, low unemployment, and stable prices. In addressing these challenges, several key elements must be considered: ① Microfoundations: According to Lucas critique, basing macroeconomic models on microeconomic behaviors is crucial. This ensures that the actions of individual agents, such as households and firms, aggregate to produce macroeconomic outcomes. ② Dynamic Interactions: Extensive interactions occur both between the macro-agent (the government) and a multitude of micro-agents (households), and within groups of households. These interactions can lead to complex economic dynamics and endogenous shocks. ③ Rational Expectations: Agents are typically modeled as rational and self-interested, making decisions based on available information and their expectations about future policies and economic conditions. ④ Stochastic Elements: Incorporating randomness and shocks, such as technological changes, policy shifts, and external economic shocks, is essential for understanding variability in economic outcomes. ⑤ Feedback Mechanisms: Feedback loops exist within the economy where macro-policy impacts micro-agents, and the micro-agents’ behaviors, in turn, influence policy decisions.
Given the aforementioned key points, Dynamic Stackelberg Mean Field Games (Dynamic SMFGs) are ideal for macroeconomic policy-making due to:
1. Leader-Follower Dynamics (②⑤): The government, acting as the leader, sets policies first. Households, as followers, adjust their behaviors based on these policies, which captures the strategic interactions between policy-setters and economic agents.
2. Mean Field Interaction (①⑤): The collective behavior of numerous micro-agents is efficiently modeled as a "mean field", which simplifies the analysis of large-scale interactions and their overall impact on the macroeconomy.
3. Dynamic Strategy (②③④): The dynamic nature of these games reflects the evolving economic conditions and policy-making, enabling the model to adapt to new information and economic shocks.
3.2 Dynamic Stackelberg Mean Field Games
@@
Following this, we consider a dynamic SMFG over discrete periods $t\in\{0,\ldots,T\}$ for macroeconomic policy-making issues: At each time $t$, the leader first selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$. Subsequently, followers simultaneously determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$. The representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$. The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
@@
Given the leader’s policy $\pi^{l}$, the dynamic SMFG is simplified into a mean field game for the followers. At time $t\in\{0,...,T-1\}$, given the leader’s action $a^{l}_{t}$, the representative follower’s state $s^{f}_{t}$ and action $a^{f}_{t}$, and the population state-action distribution of all followers denoted as $L^{f}_{t}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the representative follower receives a reward $r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and move to the next state $s_{t+1}^{f}\sim\mathbb{P}(\cdot|s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$. Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
$J^{f}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{f}\left[\sum_{t=0}^{T}r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})\right]$ (1)
where $\mathbb{E}^{f}$ denotes the expectation with respect to initial state $s_{0}^{f}\in\mu_{0}^{f}$, $(s_{t+1}^{l},s_{t+1}^{f})\sim P(\cdot\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a^{l}_{t}\sim\pi^{l}(\cdot\mid s_{t}^{l}),a^{f}_{t}\sim\pi^{f}(\cdot\mid s_{t}^{f},a^{l}_{t})$ for $t\in\{0,...,T\}$.
Definition 3.1.
[Followers’ Best Response for Leader’s Policy] Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distribution $L^{f}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L^{f})$ is defined as
${\pi^{f}}^{*}(\pi^{l},L^{f})\in\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L^{f}).$
@@
At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, and the followers’ state-action distribution $L_{t}^{f}$, the leader will receive a reward $r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})$, and move to a new state $s_{t+1}^{l}$ according to the transition probabilities $\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$. The leader aims to find the optimal policy $\pi^{l}$ to maximize the total expected reward:
$J^{l}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{l}\left[\sum_{t=1}^{T}r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})\right],$ (2)
where $\mathbb{E}^{l}$ denotes the expectation under this probability distribution for the leader when $s_{0}^{l}\sim\mu_{0}^{l},s_{t+1}^{l}\sim\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a_{t}^{l}\sim\pi_{t}^{l}(\cdot\mid s_{t}^{l})$ for $t\in\{0,...,T\}$.
Definition 3.2 (Leader’s Optimal Policy in Dynamic Stackelberg Mean Field Games).
Considering the followers’ best response $(\pi^{f},L^{f})$ for the leader’s policy, which satisfies Definition 3.1, the leader’s optimal policy ${\pi^{l}}^{*}$ in dynamic Stackelberg mean field games is defined as follows:
${\pi^{l}}^{*}$ $\in\mathop{\arg\max}_{{\pi^{l}}^{\prime}}J^{l}({\pi^{l}}^{\prime},\pi^{f},L^{f}),$

### v3

# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

@@ Abstract
The Lucas critique emphasizes the importance of considering microfoundations—how micro-agents (i.e., households) respond to policy changes—in macroeconomic policymaking. However, due to the vast scale and complex dynamics among micro-agents, predicting microfoundations is challenging. Consequently, this paper introduces a Stackelberg Mean Field Game (SMFG) approach that models macroeconomic policymaking based on microfoundations, with the government as the leader and micro-agents as dynamic followers. This approach treats large-scale micro-agents as a population, to optimize macroeconomic policies by learning the dynamic response of this micro-population. Our experimental results indicate that the SMFG approach outperforms real-world macroeconomic policies, existing AI-based and economic methods, enabling the learned macroeconomic policy to achieve the highest performance while guiding large-scale micro-agents toward maximal social welfare. Additionally, when extended to real-world scenarios, households that do not adopt the SMFG policy experience lower utility and wealth than adopters, thereby increasing the attractiveness of our policy. In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policymaking issues.
Figure 1: The mapping from macroeconomic policymaking (left) to Dynamic Stackelberg Mean Field Games (right). On the left, macro-level government actions, including fiscal or monetary policies, aim to optimize GDP, welfare, or equity. The microfoundations reflect the economic behaviors of large-scale micro-agents responding to shifts in macroeconomic policy. In Dynamic SMFGs, the government acts as the leader, and micro-agents as followers, respond to the leader’s policy.
@@ 1 Introduction
Macroeconomic policymaking is fundamental to the sustained development of an economy (Schneider and Frey 1988; Persson and Tabellini 1999). Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending (Sachs and Warner 1995). Therefore, accurately modeling and solving optimal macroeconomic policies, and evaluating the effects of their implementation are critical issues. According to the Lucas critique (Lucas Jr 1976), predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behaviors—termed “microfoundations”—to forecast micro-level responses to policy shifts and aggregate these into macroeconomic effects. Nonetheless, the vast number of micro-agents (i.e., households) and the complex dynamic interactions among them render predictions based on microfoundations exceedingly challenging.
To address the challenges outlined, this paper introduces a Stackelberg Mean Field Game (SMFG) approach: initially proposing a Dynamic Stackelberg Mean Field Game (Dynamic SMFG) framework to model macroeconomic policymaking based on microfoundations (see Figure 1). Within this framework, we design a model-free reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), aimed at optimizing macroeconomic policies by learning the dynamic responses of a large-scale population of micro-agents. In the Dynamic SMFG framework, the government acts as the leader, iteratively adjusting macroeconomic policies, while a vast array of micro-agents, serving as followers, react accordingly. This process entails sequential decision-making, with both the leader and the followers striving to optimize their cumulative long-term benefits. Given the leader’s policy, the SMFG is reduced to a mean field game, treating the large-scale micro-agents as a population to study their dynamic response to policy shifts. This dynamic SMFG framework responds to the Lucas critique. Existing models often simplify Dynamic SMFGs into single-step decisions (Pawlick and Zhu 2017), repeated games (Alcantara-Jimenez and Clempner 2020), or scenarios with myopic followers (Zhong et al. 2021), which fail to capture the complexities of macroeconomic policymaking adequately. Therefore, we introduce the SMFRL algorithm to optimize long-term utility for both the macro-agent and numerous micro-agents under the dynamic SMFG framework. In the SMFRL algorithm, we utilize mean field approximation to reduce the dimensionality of the micro-agents’ high-dimensional joint states and actions and employ a leader-follower update to learn the micro-population’s best response to macroeconomic policy adjustments, as well as to optimize macroeconomic policies considering dynamic feedback from the micro-population. This learning method enhances the stability and convergence of policies without requiring transition information or prior knowledge of the environment.
Our experiments, conducted on the TaxAI (Mi et al. 2023) environment, firstly demonstrate SMFG approach’s superiority over AI Economist (Zheng et al. ), the Saez tax (Saez 2001), the 2022 U.S. Federal tax, and free market policy (Backhouse 2005) from perspectives of macroeconomic policy evaluation and dynamic responses to economic shocks. Further, through ablation studies, we verify the significant roles of mean field approximation and leader-follower update in terms of macroeconomic policy effects, convergence, and maximal social welfare. When applied in real-world scenarios, the SMFG approach shows that households refusing the policy have lower wealth and utility than adopters, which enhances the SMFG policy’s attractiveness.
In summary, this paper’s contributions are threefold:
1. We introduce the Dynamic Stackelberg Mean Field Games framework, a novel approach for modeling macroeconomic policymaking based on microfoundations, responding to the Lucas critique.
2. We propose a model-free Stackelberg Mean Field Reinforcement Learning algorithm, designed to optimize macroeconomic policies by learning the dynamic responses of large-scale micro-agents within the Dynamic SMFG framework.
3. Experimental results demonstrate the superiority of our method in modeling and solving macroeconomic policymaking, and further explore its effectiveness in real-world settings.
Our code is available in an anonymous GitHub repository
https://anonymous.4open.science/r/SMFG˙macro˙7740
@@ 2 Related work
In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods. Classic theoretical models, such as the IS-LM model (Hicks 1980; Gali 1992), are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations. The aggregate demand-aggregate supply (AD-AS) model (Dutt 2006; Lee, Pesaran, and Smith 1997) focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics. The Solow growth model (Brock and Taylor 2010) provides a framework for economic growth but does not consider market imperfections and externalities. New Keynesian models (Blanchard and Galí 2007; Gabaix 2020) emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions. DSGE models (An and Schorfheide 2007; Smets and Wouters 2007), built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections. Empirical analysis (Ramesh et al. 2010; Vedung 2017) primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy. Econometric methods (Johnston and DiNardo 1963; Davidson, MacKinnon et al. 2004), such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality. Additionally, the Saez tax model (Saez 2001) uses earnings elasticities to optimize tax rates, enhancing social welfare without deterring economic activity. This model continues to guide current fiscal policy through effective taxation strategies.
AI for Economics
@@
Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems (Tilbury 2022). In terms of macroeconomics, AI Economist (Zheng et al. ) employs curriculum learning on tax policy design, (Trott et al. 2021) studies the collaboration between central and local governments under COVID-19, as well as monetary policy (Hinterlang and Tänzer ; Chen et al. 2023), international trade (Sch 2021), market pricing (Danassis et al. 2023). However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies. (Koster et al. ) studies democratic AI by human voting models, (Yaman et al. 2023) examines the impact of social sanction rules on labor division. However, these works are based on simplified settings and remain a gap in real-world economic policy. In terms of microeconomics, (Shi 2021; Rui and Shi 2022; Atashbar and Aruhan Shi 2023) study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium (Kuriksha 2021; Hill, Bardoscia, and Turrell 2021), multiple equilibria under real-business-cycle (Curry et al. 2022), the emergence of barter behavior (Johanson et al. 2022), optimal asset allocation and savings strategies (Ozhamaratli and Barucca 2022).
Stackelberg Mean Field Game
Stackelberg Mean Field Games (SMFGs) model the dynamic interactions between a leader and numerous homogeneous followers, enabling the development of macroeconomic policies grounded in microfoundations. The methods for SMFGs are divided into model-based and model-free algorithms. Model-based methods, such as (Bensoussan, Chau, and Yam 2015; Fu and Horst 2020; Dayanikli and Lauriere 2023; Bergault, Cardaliaguet, and Rainer 2023) employ forward-backward stochastic functional differential equations to ascertain the followers’ mean field equilibrium prior to computing the optimal strategy for the leader. Additionally, studies such as (Bensoussan et al. 2017; Lin, Jiang, and Zhang 2018; Moon and Başar 2018; Du, Wu et al. 2019; Huang and Yang 2020) simplify scenarios through linear-quadratic formulations. (Guo, Hu, and Zhang 2022), convert SMFGs into minimax optimization problems. However, these model-based methods generally depend on simplified representations of SMFGs and explicit environmental transition information. In contrast, model-free approaches leverage interaction data with the environment to learn strategies without requiring detailed knowledge of the transition dynamics and reward functions.(Pawlick and Zhu 2017) explores the resolution of SMFGs in single-step decision scenarios.(Campbell et al. 2021) utilizes a modified deep backward stochastic differential equation framework to calculate the followers’ equilibrium despite constraints on the leader’s decisions.(Miao et al. 2024) investigates locally optimal defensive strategies given predefined attacker trajectories. Furthermore,(Li et al. 2024) learns followers’ transition functions from empirical data, thus aiding the solution of SMFGs through the empirical Fokker-Planck equation. Nevertheless, the assumptions and constraints inherent in these algorithms often do not meet the complex needs of macroeconomic policymaking.
@@ 3 Stackelberg Mean Field Game Modeling
In this section, we provide details about macroeconomic policy-making problems and propose a modeling framework based on Dynamic Stackelberg Mean Field Games (Dynamic SMFGs), with the relevant mappings are shown in Figure 1.
Macroeconomic Policy-Making
To foster economic growth, enhance social welfare, or maintain social equity, governments implement economic policies such as fiscal and monetary policies. These policies directly influence micro-agents (i.e., households) and the markets they constitute, altering their expectations and behaviors. The aggregation of micro-level behaviors from numerous households manifests as macroeconomic phenomena. Therefore, in formulating policies, governments should model micro-agents’ behaviors and predict their responses to new policies, which is so-called microfoundations in the Lucas critique. However, in reality, micro-agents form a large and dynamically interacting group, each seeking to maximize personal benefits. This complexity significantly complicates predicting their responses to macroeconomic policy shifts.
To address these challenges, we propose using Dynamic SMFGs to model macroeconomic policy-making problems:
1. For microfoundations, we model the government as the leader who first sets policies, with micro-agents as followers who adjust their behaviors according to these policies. The interaction between the leader and followers forms a Stackelberg game, where the leader considers the followers’ responses to make informed decisions.
2. For large scale micro-agents, we employ a Mean Field Game to model micro-interactions among a large number of households, treating them as a population and predicting their collective response to macroeconomic policy changes.
3. For dynamic decision-making, the formulation of macroeconomic policies is inherently a multi-step decision-making process, where both the government and households make decisions at each step to optimize long-term benefits, thereby constituting a dynamic game.
Dynamic Stackelberg Mean Field Games
@@
For macroeconomic policy-making issues, we consider a dynamic SMFG with a leader and $N$ homogeneous follower agents. At any time step $t\in\{0,\ldots,T\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$. Subsequently, followers determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$. The representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$. The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively. We denote the population state-action distribution of followers under leader action $a^{l}_{t}$ and follower policy $\pi^{f}$ as $L_{t}(s^{f}_{t},a^{f}_{t};\pi^{f},a^{l}_{t})$, and abbreviate it as $L_{t}(s^{f}_{t},a^{f}_{t})$.
$L_{t}(s^{f}_{t},a^{f}_{t};\pi^{f},a^{l}_{t})\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f}),\text{where }a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$
@@
Given the leader’s policy $\pi^{l}$, the dynamic SMFG is simplified into a mean field game for the followers. At time $t\in\{0,\ldots,T-1\}$, given $(a^{l}_{t},s^{f}_{t},a^{f}_{t},L_{t})$, the representative follower then receives a reward $r^{f}(\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})$ and transitions to the next state $s^{f}_{t+1}$ according to the transition probability $s^{f}_{t+1}\sim P(\cdot|\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})$, where joint state $\mathbf{s}_{t}=\{s^{l}_{t},\mathbf{s}_{t}^{f}\}$. Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
$J^{f}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{f}\sim\mu_{0}^{f},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=0}^{T}r^{f}(\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})\right]$
where $a^{l}_{t}=\pi^{l}(s^{l}_{t})$, $a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$ and $L_{t}=L_{t}(s^{f}_{t},a^{f}_{t})$.
Definition 1 (Followers’ Best Response for Leader’s Policy).
Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distributions $L=\{L_{t}\}_{t=0}^{T}$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L)$ is defined as
${\pi^{f}}^{*}(\pi^{l},L)\in\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L).$
@@
At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, the followers’ policy $\pi^{f}$ and state-action distribution $L_{t}$, the leader will receive a reward $r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})$, and move to the next state $s_{t+1}^{l}$ according to the transition probabilities $P(\cdot|\mathbf{s}_{t},a_{t}^{l},L_{t})$. The leader aims to find the optimal policy $\pi^{l}$ to maximize the total expected reward:
$J^{l}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{l}\sim\mu_{0}^{l},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=1}^{T}r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})\right],$ (1)
where $a^{l}_{t}=\pi^{l}(s^{l}_{t})$, $a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$ and $L_{t}=L_{t}(s^{f}_{t},a^{f}_{t})$.
Definition 2 (Leader’s Optimal Policy in Dynamic Stackelberg Mean Field Games).
Considering the followers’ best response $(\pi^{f},L)$ to the leader’s policy, which satisfies Definition 1, learning the leader’s optimal policy ${\pi^{l}}^{*}$in dynamic Stackelberg mean field games is equivalent to solving the following fixed-point problem given initial condition $(\mu^{l}_{0},\mu^{f}_{0})$:
${\pi^{l}}^{*}$ $\in\mathop{\arg\max}_{{\pi^{l}}^{\prime}}J^{l}({\pi^{l}}^{\prime},\pi^{f},L)$